Assume the gold coin has a density of 19.0 g/cm3 and weighs 33.9 grams, what will be the volume of a gold coin in liters?
1
Expert's answer
2020-03-02T06:54:59-0500

density, "d= m\/v"

v=33.9/19.0=1.7842cm3 =1.7842*10-6m3

v=1.7842*10-3 liters